The escape velocity from the Earth's surface is v. The escape velocity from the surface of another planet having a radius, four times that of Earth and same mass density is :

1
Expert's answer
2021-12-27T07:55:43-0500

"v=\sqrt{gR}=\sqrt{\frac{GM}R}=\sqrt{\frac{G\rho V}R}=\sqrt{\frac{G\rho \frac 43 \pi R^3}R}=R\sqrt{\frac 43\pi \rho G,}"

"v\sim R,"

"v_2=4v_1."